# Super heavy dark matter origin of the PeV neutrino event: KM3-230213A

### v2

@@ Appendix A Alleviating the tension with IceCube’s and Auger’s null detection
As shown by Refs [1, 4], any power-law spectrum would result in a tension of $2.5-3\sigma$ between the observation of KM3-230213A by KM3NeT and null observations by IceCube and Auger. In this letter, we demonstrated that the typical nature of super heavy dark matter (SHDM) spectrum can reduce the significance of the above-mentioned tension to $<1.2\sigma$ for a DM mass $4.5\times 10^{8}$ GeV with lifetime $3.1\times 10^{29}-1.2\times 10^{30}$ s. To illustrate this, we conduct a likelihood analysis using high-energy neutrino data from ultra-high-energy neutrino observatories, KM3NeT, IceCube, and Auger, to constrain the lifetime of SHDM. Our aim is to identify DM lifetimes consistent with the observed events and to establish exclusion limits at different confidence levels based on both detections and null-detections. The number of observed events in the detectors are as follows: $N^{\rm KM3}_{\rm obs}=1$, $N^{\rm IC}_{\rm obs}=0$, and $N^{\rm Aug}_{\rm obs}=0$. We compute the expected number of events for the $i$th detector with an effective area, $A^{i}_{\rm eff}$ and an exposure time, $T^{i}_{\rm obs}$ as
[…]

### v3

@@ Appendix A Alleviating the tension with IceCube’s and Auger’s null detection
As shown by Refs [1, 4], any power-law spectrum would result in a tension of $2.5-3\sigma$ between the observation of KM3-230213A by KM3NeT and null observations by IceCube and Auger. In this letter, we demonstrated that the typical nature of super heavy dark matter (SHDM) spectrum can reduce the significance of the above-mentioned tension to $1.2\sigma$ for a DM mass $4.5\times 10^{8}$ GeV with lifetime $3.1\times 10^{29}-1.2\times 10^{30}$ s. To illustrate this, we conduct a likelihood analysis using high-energy neutrino data from ultra-high-energy neutrino observatories, KM3NeT, IceCube, and Auger, to constrain the lifetime of SHDM. Our aim is to identify DM lifetimes consistent with the observed events and to establish exclusion limits at different confidence levels based on both detections and null-detections. The number of observed events in the detectors are as follows: $N^{\rm KM3}_{\rm obs}=1$, $N^{\rm IC}_{\rm obs}=0$, and $N^{\rm Aug}_{\rm obs}=0$. We compute the expected number of events for the $i$th detector with an effective area, $A^{i}_{\rm eff}$ and an exposure time, $T^{i}_{\rm obs}$ as
$N^{i}_{\rm exp}=\int_{E_{i}}^{M_{\rm DM}/2}4\pi A^{i}_{\rm eff}(E)T^{i}_{\rm obs}\frac{d\Phi_{\nu}}{dE}dE,$ (8)